Circular No. 6115

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

COMET McNAUGHT-RUSSELL (1994u) Robert H. McNaught reports his discovery of a comet on a plate taken

by Kenneth S. Russell. The following observations are available:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Dec. 12.68490 7^{h}13^{m}15.^{s}39 −24^{o}42′42′′.8 17.5 Russell
12.71962 7 13 13.45 −24 43 04.2  _{”}
14.59184 7 11 19.85 −25 04 02.5 16.9 Kojima
14.60065 7 11 19.25 −25 04 10.2  _{”}
14.60913 7 11 18.72 −25 04 13.2  _{”}
14.67951 7 11 14.25 −25 04 57.8 16.8  _{”}

K. S. Russell (Siding Spring). U.K. Schmidt Telescope R plate. Coma diameter ∼ 0′.1 with strong central condensation and a short tail. Dense star field. Measured by McNaught.

T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6 reflector + CCD. Moonlight and bad seeing; comet low in the sky. Communicated by S. Nakano, Sumoto.

SUPERNOVA 1994ad IN ESO 152-G26 N. B. Suntzeff, Cerro Tololo Interamerican Observatory, reports that

spectrophotometry obtained on Dec. 12.2 UT with the CTIO 4-m telescope (range 320–750 nm, resolution 700) shows SN 1994ad (cf. IAUC 6103) to be a type-II supernova, very similar to the type-IIP SN 1986I (Pennypacker et al. 1989, A.J. 97, 186) six weeks after maximum.

SUPERNOVA 1994ae IN NGC 3370 Visual magnitude estimates (cf. IAUC 6111): Dec. 6.09 UT, 12.3 (T.

Vanmunster, Landen, Belgium); 11.10, 12.7 (M. Villi and G. Cortini, Mt. Colombo, Italy).

COMET MACHHOLZ (1994r) Total visual magnitude estimates (cf. IAUC 6105): Nov. 29.32 UT,

9.8 (C. S. Morris, Whitaker Peak, CA, 0.26-m reflector); Dec. 1.09, 9.8 (J. E. Bortle, Stormville, NY, 0.32-m reflector); 1.92, 9.2 (E. Broens, Mol, Belgium, 0.35-m reflector); 3.13, 9.7 (Bortle); 10.33, 10.8 (Morris).

1994 December 14 Daniel W. E. Green